# Few-shot Image Generation via Information Transfer from the Built Geodesic Surface

### 03 Jan 2024 v1

@@ Abstract
Images generated by most of generative models trained with limited data often exhibit deficiencies in either fidelity, diversity, or both. One effective solution to address the limitation is few-shot generative model adaption. However, the type of approaches typically rely on a large-scale pre-trained model, serving as a source domain, to facilitate information transfer to the target domain. In this paper, we propose a method called Information Transfer from the Built Geodesic Surface (ITBGS), which contains two module: Feature Augmentation on Geodesic Surface (FAGS); Interpolation and Regularization (I&R). With the FAGS module, a pseudo-source domain is created by projecting image features from the training dataset into the Pre-Shape Space, subsequently generating new features on the Geodesic surface. Thus, no pre-trained models is needed for the adaption process during the training of generative models with FAGS. I&R module are introduced for supervising the interpolated images and regularizing their relative distances, respectively, to further enhance the quality of generated images. Through qualitative and quantitative experiments, we demonstrate that the proposed method consistently achieves optimal or comparable results across a diverse range of semantically distinct datasets, even in extremely few-shot scenarios.
keywords:
@@ 1 Introduction
Most of the image generation methods such as Variational Auto-encoders (VAEs)kingma2013auto , Generative Adversarial Networks (GANs) goodfellow_generative_2020 , and Diffusion models ho2020denoising , have demonstrated their ability to produce images with a satisfactory combination of fidelity and diversity. Despite their impressive performance, the demanding of large-scale image datasets pose a considerable challenge for training the image generation methods. However, some certain domains, such as medical, remote sensing, and material images, pose challenges in terms of data acquisition, making it difficult to obtain the large-scale datasets typically required for training shorten2019survey ; han2023data . The scarcity of data hinders common downstream tasks like target detection, image classification, semantic segmentation, and so on. Generating images with image generation models serves as a way to effectively expand the dataset for downstream tasks. In such scenarios, the goal of image generation is to maximize the utility of a limited dataset by training a model capable of producing images that are both high in fidelity and diversity han2023data .
@@
In recent years, there have also been some studies for image generation under few-shot setting. These generative methods can be broadly categorized into two types. The first type of methods is the few-shot generative model adaption ojha_few-shot_2021 ; xiao_few_2022 , which leverages semantically relevant pre-trained image generation models in the task setup. The type of methods involves acquiring additional data to form a source domain and transferring the rich image information from the source domain to the target generator during training. It is worth noting that the scope of the adaption approach is constrained by the need for a strong semantic correlation between the source and target domain ojha_few-shot_2021 . That is to say, semantically irrelevant image information cannot be effectively transferred. Consequently, while the type of methods can be applied even in extremely few-shot scenarios, e.g, less equal than 10 samples, the source generator often necessitates numerous samples to pre-train.
@@
The second type of methods is training models from scratch without utilizing source domain karras2020training ; liu_towards_2021 ; kong_few-shot_2022 . Although these methods tend to exhibit improved performance under low-shot conditions, their effectiveness diminishes when confronted with more extreme few-shot scenarios. In such cases, the type of these methods are more susceptible to issues like overfitting and “stairlike” phenomenon kong_few-shot_2022 . Some methods, such as MixDL kong_few-shot_2022 have demonstrated the capacity to produce favorable output even in the challenging extreme few-shot scenarios, e.g., 10 samples. However, their improvement come at the expense of reduced fidelity.
@@
Finding the delicate balance between fidelity and diversity remains the top challenge in the field of extreme few-shot image generation. The type of the generative model adaption methods have exhibited noteworthy achievements in few-shot image generation methods ojha_few-shot_2021 ; xiao_few_2022 ; li_few-shot_2020 . However, acquiring an appropriate pre-trained model to serve as the source domain can be difficult in many cases. The absence of a pre-trained generator implies the absence of a readily available source domain for information transfer.
@@
To overcome the aforementioned shortcomings, we design a method called Information Transfer from the Built Geodesic Surface (ITBGS), which consists of Feature Augmentation on geodesic Surface (FAGS) module and Interpolation and Regularization (I&R) module. Figure 1 shows our motivation. FAGS focuses on the creation of a pseudo-source domain using the available few-shot training samples. Recognizing the effectiveness of data augmentation techniques in few-shot tasks qin2020diversity ; zhou2021flipda ; osahor2022ortho , we adopt the idea of data augmentation to generate the pseudo-source domain. Instead of relying on conventional data augmentation methods like horizontal or vertical flipping, brightness adjustments and color modifications, we turn to data augmentation method rooted in the Shape Space theory han2010recognition ; han2014recognizing . Instead of constructing a optimal Geodesic curve han2023fagcfeature , new data are obtained by building a Geodesic surface in the Pre-Shape Space with the training samples. Subsequently, we aim to transfer the inherent image information embedded in the pseudo-source domain to the target generator.
@@
  • 1.
    A pseudo-source domain is created for transferring the information to the target generator through adaption where no large-scale dataset or pre-trained generator exists.
  • 2.
@@
  • 3.
    Interpolation and regularization strategies are employed to enhance both image quality and diversity of interpolated images.
2 Related Work
2.1 Few-shot Image Generation
As in the conventional GAN training procedure, a dataset $\mathbb{D}_{real}$ is given to train the generator $\mathcal{G}$. The noise or latent vectors $z\sim p(z)\subset\mathbb{R}^{d}$ is drawn from the $\mathcal{D}$-dimensional latent space for better properties if using StyleGAN2 karras2019style ; zhu2020improved ; alaluf2021restyle , rather than from the Gaussian distribution. $\mathcal{G}$ maps latent vectors $z$ into generated images $\mathcal{G}(z)$ in the pixel space. The optimization function for generator $\mathcal{G}$ and a learned discriminator $\mathcal{D}$ are $L^{\mathcal{G}}_{adv}$ and $L^{\mathcal{D}}_{adv}$, respectively, defined as follows goodfellow_generative_2020 :
@@ 2.1 Few-shot Image Generation
One category of approaches is few-shot image generation with source domain, which transfer the rich image information from the source domain to the target generator, similar to transfer learning ojha_few-shot_2021 ; xiao_few_2022 ; li_few-shot_2020 . Source domain is obtained by the pre-trained image generation models, which are originated from a much larger auxiliary dataset with semantic relevance to the training set. Ojha et al. ojha_few-shot_2021 introduced a cross-domain distance consistency loss, ensuring the similarity distribution of the generated images by the target generator closely aligns with the source domain during training. The alignment aims to enhance the diversity of the generated images. Building upon the foundation, RSSA xiao_few_2022 further transferred the inherent image structure information through self-correlation consistency loss and disturbance correlation consistency loss. These loss functions effectively address potential issues such as identity degradation and image distortion.
@@
The other is few-shot image generation without source domain, which focuses on avoiding the need for additional information. Data augmentation serves as a notable solution in the context. The techniques yield impressive generation outcomes by enhancing diversity and reducing the risk of overfitting. The approaches such as DiffAugment zhao_differentiable_2020 and ADA karras2020training effectively expanded the number of real and fake samples to prevent the overfitting issue. An alternative approach involves directly altering the architecture of the generative model. For instance, a skip-layer excitation module to the generator and the discriminator is updated by a self-supervised training scheme in FastGAN liu_towards_2021 . The introduced module not only accelerates training speed and improves stability, but also facilitates the rapid and efficient generation of high-resolution images. Some methods like SinGAN shaham_singan_2019 and CoSinGAN hinz_improved_2021 include an extreme scenario, focusing on one-shot image generation. Han et al. gur2020hierarchical ; han2023data proposed an improved HP-VAE-GAN to generate material images for data augmentation. However, the improved HP-VAE-GAN lacks effectiveness for the non textured images. Most of the category of approaches struggle to achieve satisfactory output in extremely few-shot scenarios, e.g., less equal than 10 samples. Among the extreme few-shot image generation approaches, MixDL kong_few-shot_2022 stands as the single approach that utilizes no additional data or pre-trained models during the training. MixDL notably enhances the diversity of generated images, albeit with a comparatively slight compromise in fidelity.
2.2 Feature Augmentation
@@
When traversing along the feature space, it is more likely encounter realistic samples compared to the input space devries2017dataset . Feature augmentation manipulates feature vectors, rather than augments only on the image level zhang_mixup_2018 . Some methods performed simple operations on features extracted by neural networks, such as adding noise devries2017dataset and linear combination verma2019manifold . More complex transformation are also proposed for feature augmentation. For instance, a learned refinement and augmentation method is introduced in FeatMatch, which use information from prototypical class representations kuo2020featmatch . MixStyle mixed the feature statistics of two instances with a random convex weight to generate new styles zhou2021domain . Instead of directly obtaining features, Mangla et al. mangla2020charting leveraged self-supervision to obtain a suitable feature manifold before applying manifold mixup in their training procedure. Similarly, Khan et al. khan2020post generated new samples by learning a generative model over both low-level and high-level deep feature spaces. Han et al. han2023fagcfeature constructed Geodesic curve using features extracted by a pre-trained ViT, and obtained new features from the built Geodesic curves. In our method, we extract features from the continuous updating discriminator and build Geodesic surface during the every epoch of training procedure for feature augmentation.
2.3 The Shape Space Theory
@@
The Shape Space theory, originally introduced by Kendall in 1984 kendall_shape_1984 , has been a foundational concept in geometric data analysis. The Shape Space theory defines shape as the geometric information that persists when positional, scaling, and rotational effects are removed.
@@
In recent years, some interesting combination among the Shape Space theory and other domains were proposed. For instance, Kilian et al. kilian2007geometric presented continuous deformation of 3D models using Geodesic interpolation in the Shape Space. Han et al. han2010recognition ; han2014recognizing proposed a object recognition method using the Shape Space theory. They projected object contours and identified a Geodesic curve that aligns with the diverse potential shapes of the given object type in the Pre-Shape Space. Similarly, Paskin et al. paskin2022kendall projected 3D shark bone landmarks into the Shape Space and inferred the 3D pose of the shark within the 2D image on a Geodesic surface. Notably, Friji et al. friji2021geometric combined the Shape Space theory with deep learning, achieving state-of-the-art outcomes in human pose recognition tasks. However, there is currently no other work that combines the Shape Space theory with few-shot image generation.
@@
Han et al. proposed some methods to generate more new Pre-Shapes from two or three samples on the Geodesic curve or surface han2010recognition ; han2014recognizing . The Geodesic curve is derived from the following formula when provided two Pre-Shapes $\tau_{1}$ and $\tau_{2}$ in the Pre-Shape Space:
$\mathbb{G}_{curve}\left(\tau_{1},\tau_{2}\right)\left(s\right)=\left(cos(s)\right)\tau_{1}+\left(sin(s)\right)\frac{\tau_{2}-\tau_{1}cos(d({\tau_{1},\tau_{2}}))}{sin(d({\tau_{1},\tau_{2}}))},\left(0\leq s\leq d({\tau_{1},\tau_{2}})\right),$ (5)
where $d({\tau_{1},\tau_{2}})=arccos(<\tau_{1},\tau_{2}>)$, indicating the Geodesic distance between $\tau_{1}$ and $\tau_{2}$. <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > represents the dot product. The radian $s$ controls the Geodesic distance between the newly generated data point and $\tau_{1}$. By incrementing $s$ progressively, a series of gradually changing Pre-Shapes can be generated.
Given two Pre-Shapes, we can generate new Pre-Shapes using Formula 5. However, only two Pre-Shapes constrain the representation of data distribution. We expect to employ additional Pre-Shapes to generate data points that better align with the actual distribution. In cases of multiple inputs, Formula 5 becomes inapplicable. One type of approaches is to seek the optimal Geodesic curve within the Pre-Shape Space, minimizing the Geodesic distance to all input points han2010recognition ; han2023fagcfeature . The other type of approaches is to determine a Geodesic surface where all points on the surface constitute the generated Pre-Shapes. Generally, the Geodesic surface is defined in the tangent space friji2021geometric , yet errors arise from projection into the tangent space and subsequent back-projection into the Pre-Shape Space. Pennec defines the Geodesic surface as Fréchet Barycentric Subspaces 4f303e7f-6cd0-376a-9632-73d2741f0980 , which is hard to obtain data points on the Geodesic surface through the definition. We adopt the idea of some approaches that approximated the Geodesic surface with multiple Geodesic curves han2014recognizing ; paskin2022kendall .
3 Methodology
Our method, Information Transfer from the Built Geodesic surface (ITBGS), contains two module: Feature Augmentation on Geodesic Surface (FAGS); Interpolation and Regularization (I&R). We aim to strike the balance between fidelity and diversity in few-shot image generation scenarios.
3.1 Feature Augmentation on Geodesic Surface (FAGS)
@@
In fact, the information contained within the limited training samples remains underexploited and warrants further exploration. In light of this, we improve the original self-correlation consistency loss xiao_few_2022 , denoted as Geodesic self-correlation consistency loss, for the generator to capture structural information presented in the training samples. The more training samples we obtain, the richer and more accurate information they bring. Thus, for small samples, the data augmentation is especially valuable when the pre-trained models are absent. In line with existing practices, data augmentation has been widely adopted to bolster training datasets shorten2019survey . We also utilize data augmentation to create a pseudo-source domain with small samples. Subsequently, we transfer the inherent structural information from the created pseudo-source domain to the target domain. We term this approach as Feature Augmentation on Geodesic Surface (FAGS). The illustration of FAGS is shown in Figure 2.
First, we create the pseudo-source domain by performing feature augmentation, as shown in Figure 1. The features ought to be extracted from the samples to perform feature augmentation. Features extracted through GLCM haralick1973textural , SIFT lowe1999object and other deep learning based methods lecun2015deep are all suitable for feature augmentation. We focus on generating images with small samples. Thus, GAN is adopted to obtain the image features.
Most of recent studies have explored model inversion to deduce the features of input real images yin2020dreaming ; tov2021designing ; wang2022high , obtaining intermediate feature maps from the generator $\mathcal{G}$. However, model inversion necessitates a trained generator $\mathcal{G}$. In every epoch, we input real images $x\sim\mathbb{D}_{real}$ into the $l$-th layer of the discriminator $\mathcal{D}$ to extract features $\mathcal{D}^{l}(x)$ during the training procedure. The extracted features are projected into the Pre-Shape Space to build the Geodesic surface.
Generally, the Geodesic surface is defined in the tangent space friji2021geometric , yet errors arise from projection into the tangent space and subsequent back-projection into the Pre-Shape Space. Han et al. han2014recognizing approximated a Geodesic surface with multiple Geodesic curves. A Geodesic surface $\mathbb{G}_{FBS}(\tau,\omega)$ in the Pre-Shape Space can be noted as Fréchet Barycentric Subspaces (FBS) 4f303e7f-6cd0-376a-9632-73d2741f0980 , denoted as follows:
@@
$\mu$ is a vector and a Pre-Shape on the Geodesic surface $\mathbb{G}_{FBS}(\tau,\omega)$. $\tau\triangleq\{\tau_{1},...,\tau_{n}\}$ and $\omega\triangleq\{\omega_{1},...,\omega_{n}\}$, representing two sets of the given vectors and weights, respectively. $n$ denotes the number of the given vectors. Since Formula 6 is difficult to calculate, we adopt a equivalent way to accomplish the calculation with iteratively building the Geodesic curves $\mathbb{G}_{curve}(\cdot)$, from the paper paskin2022kendall . Thus, Formula 6 can be rewritten to Formula 7 based on Formula 5, as follows:
$\mu_{j}=\mathbb{G}_{curve}(\mu_{j-1},\tau_{j})(\frac{\omega_{j}}{\sum^{j}_{i=1}\omega_{i}}),\ where\ j=2,...,n,$ (7)
@@
Correspondingly, we preprocess a target domain, as shown in Figure 1. the anchor latent $\bar{z}$ kong_few-shot_2022 is calculated in the target domain using the same weights $\omega$, as follows:
@@
where $l$ iterates over the selected convolutional layers of the feature extractor. $(u,v)$ traverses all spatial positions and $L_{s\ell 1}(\cdot)$ represents the $-\ell$1 loss function ren2015faster .
3.2 Interpolation and Regularization (I&\&&R)

### 03 Mar 2024 v2

@@ Abstract
Images generated by most of generative models trained with limited data often exhibit deficiencies in either fidelity, diversity, or both. One effective solution to address the limitation is few-shot generative model adaption. However, the type of approaches typically rely on a large-scale pre-trained model, serving as a source domain, to facilitate information transfer to the target domain. In this paper, we propose a method called Information Transfer from the Built Geodesic Surface (ITBGS), which contains two modules: Feature Augmentation on Geodesic Surface (FAGS); Interpolation and Regularization (I&R). With the FAGS module, a pseudo-source domain is created by projecting image features from the training dataset into the Pre-Shape Space, subsequently generating new features on the Geodesic surface. Thus, no pre-trained models is needed for the adaption process during the training of generative models with FAGS. I&R module are introduced for supervising the interpolated images and regularizing their relative distances, respectively, to further enhance the quality of generated images. Through qualitative and quantitative experiments, we demonstrate that the proposed method consistently achieves optimal or comparable results across a diverse range of semantically distinct datasets, even in extremely few-shot scenarios.
Index Terms:
@@ I Introduction
Most of the image generation methods such as Variational Auto-encoders (VAEs)[1], Generative Adversarial Networks (GANs) [2], and Diffusion models [3], have demonstrated their ability to produce images with a satisfactory combination of fidelity and diversity. Despite their impressive performance, the demanding of large-scale image datasets pose a considerable challenge for training the image generation methods. However, some certain domains, such as medical, remote sensing, and material images, pose challenges in terms of data acquisition, making it difficult to obtain the large-scale datasets typically required for training [4, 5]. The scarcity of data hinders common downstream tasks like target detection, image classification, semantic segmentation, and so on. Generating images with image generation models serves as a way to effectively expand the dataset for downstream tasks. In such scenarios, the goal of image generation is to maximize the utility of a limited dataset by training a model capable of producing images that are both high in fidelity and diversity [5].
@@
In recent years, there have also been some studies for image generation under few-shot setting. These generative methods can be broadly categorized into two types. The first type of methods is the few-shot generative model adaption [6, 7], which leverages semantically relevant pre-trained image generation models in the task setup. The type of methods involves acquiring additional data to form a source domain and transferring the rich image information from the source domain to the target generator during training. It is worth noting that the scope of the adaption approach is constrained by the need for a strong semantic correlation between the source and target domain [6]. That is to say, semantically irrelevant image information cannot be effectively transferred. Consequently, while the type of methods can be applied even in extremely few-shot scenarios, e.g, less equal than 10 samples, the source generator often necessitates numerous samples to pre-train.
@@
The second type of methods is training models from scratch without utilizing source domain [8, 9, 10]. Although these methods tend to exhibit improved performance under low-shot conditions, their effectiveness diminishes when confronted with more extreme few-shot scenarios. In such cases, the type of these methods are more susceptible to issues like overfitting and “stairlike” phenomenon [10]. Some methods, such as MixDL [10] have demonstrated the capacity to produce favorable output even in the challenging extreme few-shot scenarios, e.g., 10 samples. However, their improvement come at the expense of reduced fidelity.
@@
Finding the delicate balance between fidelity and diversity remains the top challenge in the field of extreme few-shot image generation. The type of the generative model adaption methods have exhibited noteworthy achievements in few-shot image generation methods [6, 7, 11]. However, acquiring an appropriate pre-trained model to serve as the source domain can be difficult in many cases. The absence of a pre-trained generator implies the absence of a readily available source domain for information transfer.
@@
To overcome the aforementioned shortcomings, we design a method called Information Transfer from the Built Geodesic Surface (ITBGS), which consists of Feature Augmentation on geodesic Surface (FAGS) module and Interpolation and Regularization (I&R) module. Figure 1 shows our motivation. FAGS focuses on the creation of a pseudo-source domain using the available few-shot training samples. Recognizing the effectiveness of data augmentation techniques in few-shot tasks [12, 13, 14], we adopt the idea of data augmentation to generate the pseudo-source domain. Instead of relying on conventional data augmentation methods like horizontal or vertical flipping, brightness adjustments and color modifications, we turn to data augmentation method rooted in the Shape Space theory [15, 16]. Instead of constructing a optimal Geodesic curve [17], new data are obtained by building a Geodesic surface in the Pre-Shape Space with the training samples. Subsequently, we aim to transfer the inherent image information embedded in the pseudo-source domain to the target generator.
@@
  • A pseudo-source domain is created for transferring the information to the target generator where no large-scale dataset or pre-trained generator exists.
@@
  • Interpolation and regularization strategies are employed to enhance both image quality and diversity.
II Related Work
II-A Few-shot Image Generation
As in the conventional GAN training procedure, a dataset $\mathbb{D}_{real}$ is given to train the generator $\mathcal{G}$. The noise or latent vectors $z\sim p(z)\subset\mathbb{R}^{d}$ is drawn from the $\mathcal{D}$-dimensional latent space for better properties if using StyleGAN2 [18, 19, 20], rather than from the Gaussian distribution. $\mathcal{G}$ maps latent vectors $z$ into generated images $\mathcal{G}(z)$ in the pixel space. The optimization function for generator $\mathcal{G}$ and a learned discriminator $\mathcal{D}$ are $L^{\mathcal{G}}_{adv}$ and $L^{\mathcal{D}}_{adv}$, respectively, defined as follows [2]:
@@ II-A Few-shot Image Generation
One category of approaches is few-shot image generation with source domain, which transfer the rich image information from the source domain to the target generator, similar to transfer learning [6, 7, 11]. Source domain is obtained by the pre-trained image generation models, which are originated from a much larger auxiliary dataset with semantic relevance to the training set. Ojha et al. [6] introduced a cross-domain distance consistency loss, ensuring the similarity distribution of the generated images by the target generator closely aligns with the source domain during training. The alignment aims to enhance the diversity of the generated images. Building upon the foundation, RSSA [7] further transferred the inherent image structure information through self-correlation consistency loss and disturbance correlation consistency loss. These loss functions effectively address potential issues such as identity degradation and image distortion.
@@
The other is few-shot image generation without source domain, which focuses on avoiding the need for additional information. Data augmentation serves as a notable solution in the context. The techniques yield impressive generation outcomes by enhancing diversity and reducing the risk of overfitting. The approaches such as DiffAugment [21] and ADA [8] effectively expanded the number of real and fake samples to prevent the overfitting issue. An alternative approach involves directly altering the architecture of the generative model. For instance, a skip-layer excitation module to the generator and the discriminator is updated by a self-supervised training scheme in FastGAN [9]. The introduced module not only accelerates training speed and improves stability, but also facilitates the rapid and efficient generation of high-resolution images. Some methods like SinGAN [22] and CoSinGAN [23] include an extreme scenario, focusing on one-shot image generation. Han et al. [24, 5] proposed an improved HP-VAE-GAN to generate material images for data augmentation. However, the improved HP-VAE-GAN lacks effectiveness for the non textured images. Most of the category of approaches struggle to achieve satisfactory output in extremely few-shot scenarios, e.g., less equal than 10 samples. Among the extreme few-shot image generation approaches, MixDL [10] stands as the single approach that utilizes no additional data or pre-trained models during the training. MixDL notably enhances the diversity of generated images, albeit with a comparatively slight compromise in fidelity.
II-B Feature Augmentation
@@
When traversing along the feature space, it is more likely to encounter realistic samples compared to the input space [25]. Feature augmentation manipulates feature vectors, rather than augments only on the image level [26]. Some methods performed simple operations on features extracted by neural networks, such as adding noise [25] and linear combination [27]. More complex transformations are also proposed for feature augmentation. For instance, a learned refinement and augmentation method is introduced in FeatMatch, which use information from prototypical class representations [28]. MixStyle mixed the feature statistics of two instances with a random convex weight to generate new styles [29]. Instead of directly obtaining features, Mangla et al. [30] leveraged self-supervision to obtain a suitable feature manifold before applying manifold mixup in their training procedure. Similarly, Khan et al. [31] generated new samples by learning a generative model over both low-level and high-level deep feature spaces. Han et al. [17] constructed Geodesic curve using features extracted by a pre-trained ViT, and obtained new features from the built Geodesic curves. In our method, we extract features from the continuous updating discriminator and build Geodesic surface during the every epoch of training procedure for feature augmentation.
II-C The Shape Space Theory
@@
The Shape Space theory, originally introduced by Kendall in 1984 [32], has been a foundational concept in geometric data analysis. The Shape Space theory defines shape as the geometric information that persists when positional, scaling, and rotational effects are removed.
@@
In recent years, some interesting combination among the Shape Space theory and other domains were proposed. For instance, Kilian et al. [33] presented continuous deformation of 3D models using Geodesic interpolation in the Shape Space. Han et al. [15, 16] proposed a object recognition method using the Shape Space theory. They projected object contours and identified a Geodesic curve that aligns with the diverse potential shapes of the given object type in the Pre-Shape Space. Similarly, Paskin et al. [34] projected 3D shark bone landmarks into the Shape Space and inferred the 3D pose of the shark within the 2D image on a Geodesic surface. Notably, Friji et al. [35] combined the Shape Space theory with deep learning, achieving state-of-the-art outcomes in human pose recognition tasks. However, there is currently no other work that combines the Shape Space theory with few-shot image generation.
@@
Han et al. proposed some methods to generate more new Pre-Shapes from two or three samples on the Geodesic curve or surface [15, 16]. The Geodesic curve is derived from the following formula when provided two Pre-Shapes $\tau_{1}$ and $\tau_{2}$ in the Pre-Shape Space:
$\mathbb{G}_{cur}\left(\tau_{1},\tau_{2}\right)\left(s\right)=\left(cos(s)\right)\tau_{1}+\left(sin(s)\right)\frac{\tau_{2}-\tau_{1}cos(d({\tau_{1},\tau_{2}}))}{sin(d({\tau_{1},\tau_{2}}))},$ (5)
[…]
Given two Pre-Shapes, we can generate new Pre-Shapes using Formula 5. However, only two Pre-Shapes constrain the representation of data distribution. We expect to employ additional Pre-Shapes to generate data points that better align with the actual distribution. In cases of multiple inputs, Formula 5 becomes inapplicable. One type of approaches is to seek the optimal Geodesic curve within the Pre-Shape Space, minimizing the Geodesic distance to all input points [15, 17]. The other type of approaches is to determine a Geodesic surface where all points on the surface constitute the generated Pre-Shapes. Generally, the Geodesic surface is defined in the tangent space [35], yet errors arise from projection into the tangent space and subsequent back-projection into the Pre-Shape Space. Pennec defines the Geodesic surface as Fréchet Barycentric Subspaces [36], which is hard to obtain data points on the Geodesic surface through the definition. We adopt the idea of some approaches that approximated the Geodesic surface with multiple Geodesic curves [16, 34].
III Methodology
Our method, Information Transfer from the Built Geodesic surface (ITBGS), contains two module: Feature Augmentation on Geodesic Surface (FAGS); Interpolation and Regularization (I&R).
III-A Feature Augmentation on Geodesic Surface (FAGS)
@@
In fact, the information contained within the limited training samples remains underexploited and warrants further exploration. In light of this, we improve the original self-correlation consistency loss [7], denoted as Geodesic self-correlation consistency loss, for the generator to capture structural information presented in the training samples. The more training samples we obtain, the richer and more accurate information they bring. Thus, for small samples, the data augmentation is especially valuable when the pre-trained models are absent. In line with existing practices, data augmentation has been widely adopted to bolster training datasets [4]. We also utilize data augmentation to create a pseudo-source domain with small samples. Subsequently, we transfer the inherent structural information from the created pseudo-source domain to the target domain. We term this approach as Feature Augmentation on Geodesic Surface (FAGS). The illustration of FAGS is shown in Figure 2.
First, we create the pseudo-source domain by performing feature augmentation, as shown in Figure 1. The features ought to be extracted from the samples to perform feature augmentation. Features extracted through GLCM [37], SIFT [38] and other deep learning based methods [39] are all suitable for feature augmentation. We focus on generating images with small samples. Thus, GAN is adopted to obtain the image features.
Most of recent studies have explored model inversion to deduce the features of input real images [40, 41, 42], obtaining intermediate feature maps from the generator $\mathcal{G}$. However, model inversion necessitates a trained generator $\mathcal{G}$. In every epoch, we input real images $x\sim\mathbb{D}_{real}$ into the $l$-th layer of the discriminator $\mathcal{D}$ to extract features $\mathcal{D}^{l}(x)$ during the training procedure. The extracted features are projected into the Pre-Shape Space to build the Geodesic surface.
Generally, the Geodesic surface is defined in the tangent space [35], yet errors arise from projection into the tangent space and subsequent back-projection into the Pre-Shape Space. Han et al. [16] approximated a Geodesic surface with multiple Geodesic curves. A Geodesic surface $\mathbb{G}_{FBS}(\tau,\omega)$ in the Pre-Shape Space can be noted as Fréchet Barycentric Subspaces (FBS) [36], denoted as follows:
@@
$\mu$ is a vector and a Pre-Shape on the Geodesic surface $\mathbb{G}_{FBS}(\tau,\omega)$. $\tau\triangleq\{\tau_{1},...,\tau_{n}\}$ and $\omega\triangleq\{\omega_{1},...,\omega_{n}\}$, representing two sets of the given vectors and weights, respectively. $n$ denotes the number of the given vectors. Since Formula 6 is difficult to calculate, we adopt a equivalent way to accomplish the calculation with iteratively building the Geodesic curves $\mathbb{G}_{cur}(\cdot)$, from the paper [34]. Thus, Formula 6 can be rewritten to Formula 7 based on Formula 5, as follows:
$\mu_{j}=\mathbb{G}_{cur}(\mu_{j-1},\tau_{j})(\frac{\omega_{j}}{\sum^{j}_{i=1}\omega_{i}}),\ {\rm where}\ j=2,...,n,$ (7)
@@
Correspondingly, we preprocess a target domain, as shown in Figure 1. the anchor latent $\bar{z}$ [10] is calculated in the target domain using the same weights $\omega$, as follows:
@@
where $l$ iterates over the selected convolutional layers of the feature extractor. $(u,v)$ traverses all spatial positions and $L_{s\ell 1}(\cdot)$ represents the $-\ell$1 loss function [43].
III-B Interpolation and Regularization (I&\&&R)